Inland Revenue audits 5% of all companies every year. The companies selected for auditing in any one year are independent of the previous year’s selection. a) What is the probability that the company ‘Ross Waste Disposal’ will be selected for auditing exactly twice in the next 5 years? b) What is the probability that the company will be audited exactly twice in the next 2 years? c) What is the exact p
Answer #1
here this is binomial with parameter n=5 and p=0.05

a)

P(X=2)= (nCx)px(1−p)(n-x)    = 0.0214

(

if using ti-84 use commannd :binompdf(5,0.05,2)
if using excel use commannd :binomdist(2,5,0.05,false) )

b)

probability that the company will be audited exactly twice in the next 2 years =0.05^2 =0.0025

c) part c is not given in question)